# Online Decision-Focused Learning

### v1

@@ Abstract
Decision-focused learning (DFL) is an increasingly popular paradigm for training predictive models whose outputs are used in decision-making tasks. Instead of merely optimizing for predictive accuracy, DFL trains models to directly minimize the loss associated with downstream decisions. This end-to-end strategy holds promise for tackling complex combinatorial problems; however, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time. We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time. This setting is challenging because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex. To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) make use of the optimism principle, based on a near-optimal oracle along with an appropriate perturbation. This leads to a practical online algorithm for which we establish bounds on the expected dynamic regret, both when the decision space is a simplex and when it is a general bounded convex polytope.
1 CMAP, CNRS, École polytechnique
2 Inria Saclay, Université Paris Saclay, LMO
3 Inria Paris, Ecole Normale Supérieure, PSL
@@
Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework bertsimas2020predictive . First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making. This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management acimovic2015making ; fisher2016optimal ; ban2019big ; bertsimas2020predictive and revenue management farias2013nonparametric ; ferreira2016analytics ; cohen2017impact ; chen2022statistical to healthcare operation bertsimas2013fairness ; aswani2019data ; gupta2020maximizing ; rath2017integrated , see mivsic2020data for an extensive review. It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate. However, in practice, prediction errors are inevitable—and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
To address this limitation, an approach known as decision-focused learning mandi2024decision —also referred to as smart predict-then-optimize elmachtoub2022smart or integrated learning-optimization sadana2025survey —has emerged. Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss. By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications. While decision-focused learning yields strong empirical performance donti2017task ; verma2022case ; verma2023restless ; wang2023scalable , theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data wilder2018melding ; mandi2022decision ; shah2022decision ; schutte2024robust . This assumption breaks down in many real-world scenarios involving dynamic environments cheung2019learning ; padakandla2020reinforcement and shifting data distributions lu2018learning ; quinonero2022dataset .
Online learning cesa2006prediction ; hazan2023introductiononlineconvexoptimization provides a general way to cope with such non-stationarity of data-generating processes. This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision. Crucially, the objective function is allowed to vary over time, either in a stochastic or adversarial way. This provides a natural framework for the work that we present here, which extends decision-focused learning beyond the i.i.d., batch setting.
@@
We develop a theoretical foundation for online decision-focused learning, enabling its application in non-stationary settings. This presents a significant technical challenge, as the inherent difficulties of decision-focused learning—such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem—compound those already present in online learning. The algorithm that we develop addresses these non-differentiability and non-convexity challenges by leveraging a regularization-based approach combined with perturbation-driven optimism. More precisely, our contributions are as follows:
@@
    We present a novel algorithm—Decision-Focused Online Gradient Descent (DF-OGD) for solving this problem. Our approach relies on two main ingredients: (a) regularizing the inner problem described above to make the resulting decision differentiable, and (b) applying an original gradient-based method to the regularized outer problem, leveraging optimistic online learning rakhlin2013online together with perturbation techniques suggala2020online to handle the non-convexity of the problem.
@@
    Based on a margin assumption originating from the classification and bandit literature, we provide convergence guarantees in the form of dynamic regret bounds (defined in Section 2) for DF-OGD, when the decision space is either the simplex or a general convex polytope.
  4. (iv)
    Finally, we compare the empirical performance of DF-OGD to that of a classical prediction-focused approach in an online knapsack experiment inspired by mandi2024decision . In our simulation, our method outperforms the standard PFL approach in terms of downstream decision loss, in spite of degraded prediction loss.
@@
Our work addresses the decision-focused learning problem, pioneered by elmachtoub2022smart . Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions gould2016differentiating ; amos2017optnet ; donti2017task ; wilder2018melding ; mandi2020interior , smoothing via random perturbation berthet2020learning , building surrogate losses via duality elmachtoub2022smart or directional gradients huang2024decision and relying on pairwise ranking techniques mandi2022decision , see mandi2024decision ; sadana2025survey for additional references. However, the extension of decision-focused learning to the online setting remains unexplored. While several recent studies address online bi-level optimization shen2023online ; tarzanagh2024online ; lin2023non , none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems. In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
This motivates the use of different online methods, specifically tailored to address these problems. On the one hand, lack of differentiability is usually tackled through zero-th order methods heliou2020gradient ; frezat2023gradient , sub-gradient duchi2011adaptive , proximal dixit2019online or smoothing approaches abernethy2014online . On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle kalai2005efficient ; agarwal2019learning ; suggala2020online ; xu2024online or additional smoothness conditions lesage2020second ; ghai2022non . All those results are derived for static regret zinkevich2003online ; hazan2023introductiononlineconvexoptimization which compares the learner efficiency to the best static strategy (see Appendix A). An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret hazan2017efficient ; aydore2019dynamic ; zhuang2020no ; hallak2021regret , which is the sum of the objective gradient norms evaluated in the iterates over time. Minimizing local regret corresponds to encouraging convergence toward stationary points. However as we shall see, this approach is not sensible in our framework, as objective gradients are either zero or undefined. In this work, we focus on dynamic regret, defined in Section 2. Originally introduced in zinkevich2003online and later developed in hall2013dynamical ; besbes2015non ; zhao2020dyn ; zhao2021improved among others, this regret takes into account the evolution of the environment and is well suited to our framework. We also provide an analysis for the more classic static regret notion based on techniques from suggala2020online , see Appendix B.
@@
In Section 2, we introduce the online decision-focused problem, our regret notion and our assumptions. In Section 3, we present the DF-OGD algorithm, before analyzing its dynamic regret performance in the simplex and convex polytope cases in Section 4.1 and Section 4.2 respectively. Finally, we present our experiment in Section 5.
@@
For a differentiable map $g:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla g(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}g(x)$ denotes the Hessian of $g$. In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla g(x)$ denotes the Jacobian of $g$. For two vectors $(v,w)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $v\succcurlyeq w$ means that $v_{i}\geqslant w_{i}$ for any $i\in[d]$, and $\mathopen{}\mathclose{{}\left\langle v,w}\right\rangle=v^{\mathsf{\scriptscriptstyle T}}w$ refers to the standard Euclidian inner product. Also, $\mathopen{}\mathclose{{}\left\|v}\right\|=\sqrt{\mathopen{}\mathclose{{}\left\langle v,v}\right\rangle}$ is the standard euclidian norm. Given a compact convex set $\Theta\subseteq\mathbb{R}^{d}$, $\Pi_{\Theta}$ denotes the orthogonal projection onto $\Theta$. For a matrix $M\in\mathbb{R}^{m\times d}$, $\mathopen{}\mathclose{{}\left\|M}\right\|_{\mathrm{op}}$ refers to its $\mathrm{L}^{2}$-operator norm. In the case $d=m$, $\lambda_{\min}(M)$ refers to its lowest real eigenvalue and $\lambda_{\max}(M)$ it largest real eigenvalue. For $(x,y)\in\mathbb{R}^{2}$, $x\propto y$ means that there exists $\lambda\in\mathbb{R}^{\star}$ such that $x=\lambda y$.
@@ 2 Framework
where $X_{t}\in\mathcal{X}$ are covariates (possibly random), $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$, and $\bar{g}_{t}:\mathcal{X}\to\mathcal{Z}$, $\mathcal{Z}\subset\mathbb{R}^{d}$ is a cost function, which for instance satisfies $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathcal{Z}\subset\mathbb{R}^{d}$. Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round. While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathcal{Z}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it. Then, the decision-maker seeks to jointly fit $\theta$ and solve the resulting decision problem. More precisely, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion. At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$. Formally, they make the decision $w_{t}=w^{\star}_{t}(\theta_{t})$ where for any $\theta\in\Theta$,
$w_{t}^{\star}(\theta)\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\langle g(\theta,X_{t}),w\rangle\;.$ (2)
Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$. Finally, they update their prediction parameter $\theta_{t+1}$ based on $\mathcal{H}_{t}$ through an algorithm $\mathsf{Alg}_{t}$. A central question for the decision-maker is then the choice of algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$ to fit their regression model, that is how to pick $\theta_{t}\in\Theta$ for each $t\in[T]$.
@@
From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a risk function based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as an empirical risk of the form $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g. cross-entropy or squared error).
Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function. Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2). From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem colson2007overview ; sinha2017review ; ji2021bilevel .
We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$, allowing an adversarial selection of the features $X_{t}$ and cost functions $\bar{g}_{t}$. To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with a dynamic regret. It compares the learner’s performance against a sequence of best decisions that may vary over time. More precisely, let
@@
denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$. In line with zinkevich2003online ; haddouche2024optimisticallytemperedonlinelearning , we define the following dynamic regret:
$\mathfrak{R}_{T}=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}F_{t}(\theta)\;,\quad\text{where}\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{}\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}\right]\;.$ (4)
We also consider a static regret and derive associated bounds in Appendix B. We make the following mild regularity assumptions for the rest of the analysis.
@@
(i) $\Theta$ is a compact and convex set with diameter $D_{\Theta}<\infty$. (ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})$ is continuously differentiable and $\mathopen{}\mathclose{{}\left\|\nabla_{\theta}g(\theta,X_{t})}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely. (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space boyd2004convex ; bishop2006pattern . Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter. We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning bennouna2024addressing .
Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard. Therefore, we make an assumption coming from the classification mammen1999smooth ; tsybakov2004optimal and bandit literature zeevi2009woodroofe ; perchet2013multi about the margins of the cost function. Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$. We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{}\left\langle g(\theta,x),v_{i}}\right\rangle$ for $i\in[d]$, and assume that
@@
There exists $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{}\left[0,1}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{}\left[0,1}\right]$,
@@
H 2 controls how difficult it is to solve problem (2) since it determines the objective gap between the optimal vertex $v_{I_{t}(\theta)}\in\mathcal{W}$ and the other vertices. In other words, it quantifies how identifiable is the optimal vertex, $\varepsilon=0$ meaning it is not distinguishable from the others. Remark that H 2 is only related to the distribution of the features $\{X_{t}\}_{t\in[T]}$. If it is satisfied for $\beta\geqslant 1$, then it is automatically satisfied for $\beta=1$ and therefore we do not loose any generality restraining $\beta$ to $\mathopen{}\mathclose{{}\left[0,1}\right]$. This assumption is critical in our analysis, as it allows to bound the expected distance between the actual optimal decision $w^{\star}_{t}(\theta)$ and the regularized approximation $\tilde{w}_{t}(\theta)$ introduced in (5) below.
3 An online algorithm for decision-focused learning.
A natural approach to minimize the regret (4) would be to use an Online Gradient Descent (OGD) scheme zinkevich2003online for the algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$, which is standard in online learning. However, since in (2), $w^{\star}_{t}(\theta)$ minimizes a linear function over the convex polytope $\mathcal{W}$, $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$, thus $\nabla w^{\star}_{t}(\theta)$ is either zero or undefined.
To address this issue, following wilder2018melding , we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,

### v2

@@ Abstract
Decision-focused learning (DFL) is an increasingly popular paradigm for training predictive models whose outputs are used in decision-making tasks. Instead of merely optimizing for predictive accuracy, DFL trains models to directly minimize the loss associated with downstream decisions. However, existing studies focus solely on scenarios where a fixed batch of data is available and the objective function does not change over time. We instead investigate DFL in dynamic environments where the objective function and data distribution evolve over time. This setting is challenging for online learning because the objective function has zero or undefined gradients—which prevents the use of standard first-order optimization methods—and is generally non-convex. To address these difficulties, we (i) regularize the objective to make it differentiable and (ii) use perturbation techniques along with a near-optimal oracle to overcome non-convexity. Combining those techniques yields two original online algorithms tailored for DFL, for which we establish respectively static and dynamic regret bounds. These are the first provable guarantees for the online decision-focused problem. Finally, we showcase the effectiveness of our algorithms on a knapsack experiment, where they outperform two standard benchmarks.
@@
Many real-world decision problems involve uncertainty, and a common approach to handling it is through the predict-then-optimize framework [10]. First, a prediction model is trained on historical data; then, its output is fed into an optimization problem to guide decision-making. This natural strategy has been successfully applied in many operation research (OR) problems, ranging from supply chain management [2, 28, 7, 10] and revenue management [26, 27, 18, 16] to healthcare operation [11, 5, 34, 61], see Mivsić and Perakis [56] for an extensive review. It is clear that this approach would yield optimal decisions if the predictions were perfectly accurate. However, in practice, prediction errors are inevitable—and even small inaccuracies can propagate through the optimization process, potentially leading to poor decisions.
To address this limitation, an approach known as decision-focused learning [54]—also referred to as smart predict-then-optimize [25] or integrated learning-optimization [62]—has emerged. Instead of optimizing for prediction accuracy alone, this method trains the predictive model to directly minimize the downstream decision loss. By aligning the learning objective with the decision-making goal, it produces models that are more robust to prediction errors in practical applications. While decision-focused learning yields strong empirical performance [23, 71, 72, 73], theoretical development has so far been limited to the batch setting, where models are trained on pre-collected, independently and identically distributed (i.i.d.) data [74, 53, 64, 63]. This assumption breaks down in many real-world scenarios involving dynamic environments [17, 57] and shifting data distributions [50, 60].
Online learning [15, 37] provides a general way to cope with such non-stationarity of data-generating processes. This framework considers a learner who makes decisions sequentially, at each round leveraging data collected from previous rounds to inform its next decision. Crucially, the objective function is allowed to vary over time, either in a stochastic or adversarial way. This provides a natural framework for the work that we present here, which extends decision-focused learning beyond the i.i.d., batch setting.
@@
We develop a theoretical foundation for online decision-focused learning, enabling its application in non-stationary settings. This presents a significant technical challenge, as the inherent difficulties of decision-focused learning—such as the non-differentiability of the objective function or the lack of convexity of the losses due to the bi-level nature of the problem—compound those already present in online learning. Our contributions are as follows:
@@
    We present two algorithms to tackle this problem, Decision-Focused Follow-the-Perturbed-Leader (DF-FTPL) and Decision-Focused Online Gradient Descent (DF-OGD). While both rely on regularizing the inner problem to make the resulting decision differentiable, they differ on the way they update the parameters of the prediction model. DF-FTPL uses the FTPL approach [43], and DF-OGD leverages a variant of Online Gradient Descent [81]. We establish sublinear convergence guarantees for both procedures, in the form of a static regret bound for the former and a dynamic regret bound for the latter. To our knowledge, these are the first provable guarantees for the online decision-focused learning problem.
@@
    Finally, we assess the performance of our algorithms on a knapsack experiment inspired by [54]. Our simulations demonstrate that our approach outperforms the online version of two popular baselines, namely prediction-focused learning and Smart-Predict-then-Optimize, in both static and dynamic environments.
@@
Our work addresses the decision-focused learning problem, pioneered by [25]. Several approaches have been proposed in the batch setting, among which differentiating the associated KKT conditions [33, 4, 23, 74, 52], smoothing via random perturbation [9], building surrogate losses via duality [25] or directional gradients [42] and relying on pairwise ranking techniques [53], see [54, 62] for additional references. However, the extension of decision-focused learning to the online setting remains unexplored. While several recent studies address online bi-level optimization [65, 69, 49], none of them are applicable to decision-focused learning as they rely on restrictive smoothness assumptions, that are incompatible with the structure of decision-focused problems. In particular, the decision-focused objective is typically non-convex and features gradients that are either zero or undefined, owing to its underlying linear structure.
This motivates the use of different online methods, specifically tailored to address these problems. On the one hand, lack of differentiability is usually tackled through zero-th order methods [39, 29], sub-gradient [24], proximal [22] or smoothing approaches [1]. On the other hand, methods to address non-convexity in online learning often rely on the existence of a near optimal oracle [45, 3, 67, 75] or additional smoothness conditions [46, 32]. All those results are derived for static regret [81] which compares the learned predictors to the best static strategy and is defined in Section˜2. A more challenging criterion, also described in Section˜2 is dynamic regret, introduced in [81] and later developed in [35, 12, 79, 78] among others, which takes into account the evolution of the environment. An alternative approach to obtain guarantees in the online non-convex setting is to weaken static regret to local regret [38, 6, 80, 36], which is the sum of the objective gradient norms evaluated in the iterates over time. Minimizing local regret corresponds to encouraging convergence toward stationary points. However as we shall see, this approach is not sensible in our framework, as objective gradients are either zero or undefined.
@@
In Section˜2, we introduce the online decision-focused problem, our notions of regret and our assumptions. In Section˜3, we present the DF-FTPL and DF-OGD algorithms, before deriving bounds on the static regret of the former and the dynamic regret of the latter. Finally, we present our experiment in Section˜4.
@@
For a differentiable map $\varphi:\mathbb{R}^{m}\rightarrow\mathcal{Y}$ such that $\mathcal{Y}\subseteq\mathbb{R}$, $\nabla\varphi(x)$ denotes the gradient of $g$ at $x\in\mathbb{R}^{m}$ and $\nabla^{2}\varphi(x)$ denotes the Hessian of $g$. In the case where $\mathcal{Y}\subseteq\mathbb{R}^{d}$, $\nabla\varphi(x)$ denotes the Jacobian of $\varphi$. For two vectors $(v,w)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, $v\succcurlyeq w$ means that $v_{i}\geqslant w_{i}$ for any $i\in[d]$, and $\mathopen{}\mathclose{{\left\langle v,w}}\right\rangle=v^{\mathsf{\scriptscriptstyle T}}w$ refers to the standard Euclidian inner product. Also, $\mathopen{}\mathclose{{\left\|v}}\right\|=\sqrt{\mathopen{}\mathclose{{\left\langle v,v}}\right\rangle}$ is the standard euclidian norm. Given a compact convex set $\Theta\subseteq\mathbb{R}^{d}$, $\Pi_{\Theta}$ denotes the orthogonal projection onto $\Theta$. For a matrix $M\in\mathbb{R}^{m\times d}$, $\mathopen{}\mathclose{{\left\|M}}\right\|_{\mathrm{op}}$ refers to its $\mathrm{L}^{2}$-operator norm. In the case $d=m$, $\lambda_{\min}(M)$ refers to its lowest real eigenvalue and $\lambda_{\max}(M)$ it largest real eigenvalue. For $(x,y)\in\mathbb{R}^{2}$, $x\propto y$ means that there exists $\lambda\in\mathbb{R}^{\star}$ such that $x=\lambda y$.
@@ 2 Framework
where $X_{t}\in\mathcal{X}$ are covariates (possibly random) and $\mathcal{W}=\mathrm{Conv}(v_{1},\ldots,v_{K})$ is a bounded convex polytope of $\mathbb{R}^{d}$ with non-empty interior and vertices $\{v_{i}\}_{i=1}^{K}$. This feasible set appears naturally in many problems such as shortest-path [30], portfolio selection [47] or mixed strategy design in games [68]. In (1), $\bar{g}_{t}:\mathcal{X}\to\mathbb{R}^{d}$, is a cost function, which satisfies for instance $\bar{g}_{t}(X_{t})=\mathbb{E}[Z_{t}\mid X_{t}]$ for some hidden state $Z_{t}\in\mathbb{R}^{d}$. Importantly, $X_{t}$ is revealed at the beginning of the round, but not $\bar{g}_{t}(X_{t})$ which is only available at the end of the round. While $\bar{g}_{t}(X_{t})$ is unknown to the decision-maker at the decision time, they have access to a family of models $g:\Theta\times\mathcal{X}\rightarrow\mathbb{R}^{d}$, parameterized by $\Theta\subset\mathbb{R}^{m}$, to predict it. Then, the general form of the decision-making dynamics we consider can be described as follows: at each round $t\in[T]$, for a horizon $T\in\mathbb{N}$, given the current parameter $\theta_{t}$,
  1. 1.
    The decision-maker observes $X_{t}$ and compute its prediction as $g(\theta_{t},X_{t})$.
  2. 2.
    Then, they take an action minimizing the resulting predicted cost
    $w_{t}=w^{\star}_{t}(\theta_{t})\in\operatorname*{\arg\!\min}_{w\in\mathcal{W}}\mathopen{}\mathclose{{\left\langle g(\theta_{t},X_{t}),w}}\right\rangle$ (2)
  3. 3.
    Finally, they observe $\bar{g}_{t}(X_{t})$ and update $\theta_{t+1}$ for the next round.
Formally, the decision-maker considers a joint process $(\theta_{t},w_{t})_{t\in[T]}$, starting from $\theta_{1}\in\Theta$ and a history $\mathcal{H}_{0}=\emptyset$, and defined by the following recursion. At round $t\in[T]$, the decision-maker takes the best action given the collected history $\mathcal{H}_{t-1}$, the current estimate $\theta_{t}$ and a new feature $X_{t}$ as in (2). Then, they observe $\bar{g}_{t}(X_{t})$ and update their history $\mathcal{H}_{t}=\mathcal{H}_{t-1}\cup\{(X_{t},\bar{g}_{t}(X_{t}),\theta_{t},w^{\star}_{t}(\theta_{t}))\}$. Finally, they update their prediction parameter $\theta_{t+1}$ based on $\mathcal{H}_{t}$ through an algorithm $\mathsf{Alg}_{t}$. A central question for the decision-maker is then the choice of algorithms $\{\mathsf{Alg}_{t}\}_{t\in[T]}$ to fit their regression model, that is how to pick $\theta_{t}\in\Theta$ for each $t\in[T]$.
@@
From an online perspective, prediction-focused learning consists in selecting at each $t\in[T]$ an algorithm $\mathsf{Alg}_{t}$ to estimate $\operatorname*{\arg\!\min}_{\theta\in\Theta}\mathrm{R}_{t}(\theta)$, where $\mathrm{R}_{t}$ is a statistical risk based on the historical observations $\mathcal{H}_{t-1}$, typically chosen as the empirical risk $\mathrm{R}_{t}(\theta)=\sum_{i=1}^{t}\ell(\bar{g}_{i}(X_{i}),g(\theta,X_{i}))$, for some loss function $\ell$ (e.g., cross-entropy or squared error).
Alternatively, we consider in this paper, the decision-focused learning approach, which selects parameters $\theta\in\Theta$ by directly minimizing the downstream objective instead of a risk function. Specifically, the decision-maker chooses $\mathsf{Alg}_{t}$ to solve $\operatorname*{\arg\!\min}_{\theta\in\Theta}\langle\bar{g}_{t}(X_{t}),w_{t}^{\star}(\theta)\rangle$ where $w_{t}^{\star}$ is defined in (2). From the previous formulation, we see that the decision-focused learning formulation corresponds to a bilevel optimization problem [19, 66, 44].
We emphasize that no stationary assumptions are made on the process $\{X_{t},\bar{g}_{t}(X_{t})\}_{t\in[T]}$, allowing an adversarial selection of the features $X_{t}$ and cost functions $\bar{g}_{t}$. To fully accommodate this flexibility, we assess the optimality of $\{\theta_{t}\}_{t\in[T]}$ with two notions of regret. Let
@@
denotes the loss incurred when taking an action based on the prediction parameter $\theta_{t}\in\Theta$. Inspired from [81], we consider the notions of static and dynamic regret to measure the effectiveness of a learning strategy $\mathopen{}\mathclose{{\left\{\theta_{t}}}\right\}_{t\in[T]}$, which are respectively:
$\mathfrak{R}^{s}_{T}=\sum_{t\in[T]}f_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)}}\right]\quad\text{and}\quad\mathfrak{R}^{d}_{T}=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}F_{t}(\theta),$ (4)
where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$. Note that our static regret compare to the best fixed strategy, with respect to the averaged losses. This notion makes sense here as we aim to control those regrets in expectation over the randomness of the process. In our dynamic regret, we consider the expectation of the loss given past information to prevent the dynamic comparator from overfitting the current realization of $X_{t}$.
We make the following mild regularity assumptions for the rest of the analysis.
@@
(i) $\Theta\subset\mathbb{R}^{m}$ is a compact and convex set with diameter $D_{\Theta}<\infty$. (ii) For any $\theta\in\Theta$ and $t\in[T]$, $\nabla_{\theta}g(\theta,X_{t})\in\mathbb{R}^{m\times d}$ is continuously differentiable and $\mathopen{}\mathclose{{\left\|\nabla_{\theta}g(\theta,X_{t})}}\right\|_{\mathrm{op}}\leqslant G<\infty$ almost-surely. (iii) For any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|\leqslant D_{\mathcal{Z}}<\infty$ almost-surely.
It is common in online learning to assume boundedness of the parameter space, model gradient and prediction space [14, 13]. Note that in the well-specified setting, i.e., $\bar{g}_{t}=g(\bar{\theta}_{t},\cdot)$ for some $\bar{\theta}_{t}$, these two former conditions automatically imply the latter. We emphasize however that we do not assume this realizability condition, in contrast to most of the literature on decision-focused learning [8].
Without further assumptions on the sequence of costs $\bar{g}_{t}$ and models $g$, the problem can still be made arbitrarily hard. Therefore, we make an assumption coming from the classification [51, 70] and bandit literature [76, 59] about the margins of the cost function. Recall that we denote by $\{v_{i}\}_{i=1}^{K}$ the vertices of $\mathcal{W}$. We define for any $x\in\mathcal{X}$, $u_{i}(\theta,x)=\mathopen{}\mathclose{{\left\langle g(\theta,x),v_{i}}}\right\rangle$ for $i\in[d]$.
@@
There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{\left[0,1}}\right]$, such that almost surely for any $t\in[T]$, $\theta\in\Theta$ and $\varepsilon\in\mathopen{}\mathclose{{\left[0,1}}\right]$,
@@
H˜2 controls how difficult it is to solve problem (2) since it determines the objective gap between the optimal vertex $v_{I_{t}(\theta)}\in\mathcal{W}$ and the other vertices. In other words, it quantifies how identifiable is the optimal vertex, $\varepsilon=0$ meaning it is not distinguishable from the others. If it is satisfied for $\beta\geqslant 1$, then it is automatically satisfied for $\beta=1$ thus we do not loose any generality restraining $\beta$ to $\mathopen{}\mathclose{{\left[0,1}}\right]$. This assumption is critical in our analysis, as it allows bounding the expected distance between the actual optimal decision $w^{\star}_{t}(\theta)$ and the regularized approximation $\tilde{w}_{t}(\theta)$ introduced in (5) below.
We expect H˜2 to hold in a wide variety of classical statistical settings. For example, in Appendix˜A we show that it is satisfied when $g:(\theta,X)\mapsto X\theta$ where $X$ has i.i.d standard Gaussian columns.
3 Online algorithms for decision-focused learning .
On the need of regularization to get differentiation.
From an online learning perspective, a natural approach to minimize the two regrets defined in (4) would be to update $\theta_{t+1}$ at each round $t$ using a variant of either the Follow-The-Leader algorithm [FTL, 45] or Online Gradient Descent [OGD, 81], applied to the objective function $f_{t}$ defined in (3), thus specifying the choice of $\mathsf{Alg}_{t}$. However, these algorithms are designed for single-level optimization problems and require access to (sub-)gradients of the objective. In our setting, this requirement is problematic because the function $f_{t}$ does not admit an informative gradient: by construction, computing the gradient of $f_{t}$ involves differentiating the mapping $\theta\mapsto w_{t}^{\star}(\theta)$, which is generally not (sub-)differentiable. Indeed, $w^{\star}_{t}$ minimizes a linear function over the convex polytope $\mathcal{W}$, thus $\theta\mapsto w_{t}^{\star}(\theta)$ ranges in the finite set of vertices $\{v_{1},\ldots,v_{K}\}$.
To address this issue, following [74], we propose to add a regularizer $\mathcal{R}$ to the objective function in (2). Accordingly, we define for any $\theta\in\Theta$,